OpenStudy (anonymous):

sqrt{4}{256} sqrt{4}{256} @Mathematics

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[\sqrt[4]{256} \]

OpenStudy (anonymous):

all that trouble just to make sure i had it correct lolol oh me

OpenStudy (anonymous):

u are making mistake for click on equation then in the options go for sqrt then type

OpenStudy (jamesj):

Yep. \[ (256)^{1/4} = (2^8)^{1/4} = 2^{8/4} = 2^2 = 4 \]

OpenStudy (anonymous):

how come i cant get it to show up like that in the room? clicking equation is the only way i know how to do it, either that or click draw

OpenStudy (anonymous):

no u can type it in equation because from that only question is easy to understand

OpenStudy (anonymous):

i guess if anyone replies to it i can just redo it here